OpenStudy (anonymous):

Which of the following is true? The base of a logarithmic function can be a negative number. A logarithmic function is the inverse of a rational function. The domain of the logarithmic function f(x) = log5x is all positive real numbers. The base of a logarithmic function can be zero.

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

First one

jimthompson5910 (jim_thompson5910):

The base of a logarithmic function can be a negative number. False A logarithmic function is the inverse of a rational function. False The domain of the logarithmic function f(x) = log5x is all positive real numbers. True The base of a logarithmic function can be zero. False

jimthompson5910 (jim_thompson5910):

It's not the first since you can't take the log of a negative number

OpenStudy (anonymous):

none

OpenStudy (anonymous):

yeah , but this asking for base of log

jimthompson5910 (jim_thompson5910):

use the change of base formula and you'll see it

OpenStudy (anonymous):

Base has to be positive...

OpenStudy (anonymous):

\[\log_{-2}x=b\] \[(-2)^b=x\]

OpenStudy (anonymous):

"The logarithm is denoted "logb(y)" (pronounced as "the logarithm of y to base b" or "the "base-b logarithm of y"). In the equation logb(y) = x, the value x, is the answer to the question "To what power must b be raised, in order to yield y?". For the logarithm to be defined, the base b must be a positive real number not equal to 1 and y must be a positive number." -Wikipedia

jimthompson5910 (jim_thompson5910):

\[\log_{-2}(x)=\frac{\log(x)}{\log(-2)}\]

OpenStudy (anonymous):

oh, I see,you are right

OpenStudy (anonymous):

The domain of the logarithmic function f(x) = log5x is also false because there isnt such a thing as log (0) !!!

jimthompson5910 (jim_thompson5910):

The domain of the logarithmic function f(x) = log5x is all POSITIVE real numbers is true

OpenStudy (anonymous):

It's the third one....

OpenStudy (anonymous):

a positive number is >= to 0

jimthompson5910 (jim_thompson5910):

no, 0 is not a positive number nor is it negative

OpenStudy (anonymous):

a strictly positive number > 0

OpenStudy (anonymous):

That's silly. It has nothing to do with the change of base formula. It has to do with the fact that (-2)^x is not a one to one function and doesn't have an inverse.

OpenStudy (anonymous):

Or any negative base for that matter.

jimthompson5910 (jim_thompson5910):

well he was pointing out an example of a log with a negative base, which can be shown is not possible using the change of base formula

OpenStudy (shamil98):

Yeah, its da third one.